Rob is considering investing in a new method of cleaning carpets that will cost him more both in fixed and variable costs, but will allow him to charge more. He wants to know what his total sales will be for both his current method and the new method at his breakeven point. You will need to calculate his BEv for both methods first and then determine his total sales for BOTH to give him the information he needs. Be sure you round UP to the nearest whole carpet. Remember the current has these factors: Price to clean a carpet = $75 Materials (variable costs) = $15 Monthly fixed costs = $4000 The new method has these factors: Fixed costs: $4750 Variable costs per carpet: 18.50 Price per carpet: $99 What are Rob’s sales for the current method AND the new method, in that order? Question 5 options: 1) $5010 and $5841 2) $5025 and $5950 3) $5025 and $5841 4) $5005 and $5775 5) There is not enough information to determine this
Answer #1

Answer - 3) $5025 and $5841

BEv = Fixed cost/ (Price – Unit Variable Cost)
Current method
BEv = 4000 / (75 – 15) = 67 carpets
New method
BEv = 4750 / (99 – 18.50) = 59 carpets


Sales = BEv X Unit price
Current method
Sales = 67 x 75 = $5025
New method
Sales = 59 x 99 = $5841